# When LLMs Meet Cunning Questions: A Fallacy Understanding Benchmark for Large Language Models

### v1

@@ 3.1 Experimental Settings
To better reflect the evaluation of FLUB’s fallacy understanding ability of LLMs, we select some advanced LLMs that are widely used and have great influence in the Chinese community:
  • ERNIE-Bot Baidu (2023) is a series of closed-sourced commercial LLMs released by Baidu. We evaluate the three latest chat models in the ERNIE-bot series, including ERNIE-Bot-3.5, ERNIE-Bot-3.5-Turbo, and ERNIE-Bot-4.0.
  • ChatGPT OpenAI (2023) ChatGPT is undoubtedly the hottest and most studied model developed by OpenAI. Currently, ChatGPT mainly has GPT-3.5 and GPT-4, so we evaluate GPT-3.5-Turbo and GPT-4-Turbo.
  • ChatGLM3 Du et al. (2022) is the latest open-sourced model of the ChatGLM series, and ChatGLM is a series of bilingual LLMs based on the General Language Model (GLM) framework. We evaluate the only open-sourced parameter size of ChatGLM3, which is 6B (i.e., ChatGLM3-6B).
  • Qwen Bai et al. (2023) is a series of open-sourced LLMs that encompasses distinct models developed by the Alibaba Group. We select three chat Qwen models with various parameter sizes, including Qwen-7B-Chat, Qwen-14B-Chat, and Qwen-72B-Chat.
  • Yi 01-AI (2023) series models are open-sourced LLMs trained from scratch by 01-AI. The Yi models are trained on large-scale multilingual corpus as the bilingual models. In our experiments, we select Yi-6B-Chat and Yi-34B-Chat to be evaluated on FLUB.
  • Baichuan2 Yang et al. (2023) is a series of open-sourced multilingual models that have achieved the competitive performance of its size on many Chinese benchmarks. Based on their open-sourced status, Baichuan2-7B-Chat and Baichuan2-13B-Chat are selected by us.
In the process of our running LLMs inference, for closed-sourced LLMs, we access corresponding models via the official APIs. Meanwhile, open-sourced models are deployed on 1 to 4 NVIDIA A100 GPUs depending on their parameter size.
@@ 3.2 Automatic Evaluation Results
For Task 1 and Task 2, we automatically calculate the accuracy. For Task 3, we utilize GPT-4 to automatically score the explanations generated by LLMs. The automatic results are presented in Table 2 and we have the following insights:
@@
    For the difficulty of different tasks, as we expected, the Answer Selection task is the simplest, which shows that LLMs should have a certain ability to distinguish right from wrong when seeing correct and wrong answers. However, we also see that the performance of all models on the Question Type Classification task is unsatisfactory, with accuracy rates below 25%. This deficiency may stem from the models’ limited capability to comprehend the semantics of various question categories.
@@
    For the model performance of different scale parameters, overall, models of larger scale are better equipped to understand cunning questions, which aligns with intuitive expectations. Of course, there are exceptions. We find that Qwen-7B-Chat achieves better performance on the Question Type Classification task than other larger models, which indicates that Qwen-7B-Chat’s fallacy type understanding ability is very competitive.
@@
    For the connection between different tasks, the comparative outcomes among different models across the three tasks are not consistent. Nevertheless, models that exhibit superior performance in the Answer Selection task tend to generate more plausible explanations. This phenomenon reminds us that there is a close relationship between the Answer Selection task and the Question Explanation task. The interaction between these two tasks is very critical for improving the fallacy understanding ability of LLMs.
@@
    For the impact of Chain-of-Thought, to our surprise, judging from the results, Chain-of-Thought does not bring qualitative improvements to LLMs’ reasoning ability and fallacy understanding ability. Especially for the Answer Selection task, Chain-of-Thought even has negative impacts. This phenomenon demonstrates the challenging nature of FLUB and implies that we need to study new strategies besides Chain-of-Thought to stimulate LLMs’ reasoning capabilities.
@@ 3.3 The Impact of In-context Learning
We select 5 high-performing LLMs to study the impact of in-context learning on LLMs’ fallacy understanding ability. Demonstrations used for in-context learning are randomly selected. As shown in Figure 4, unlike Chain-of-Thought which has almost no positive effect, the LLM’s performance with in-context learning is basically on the rise as demonstrations increase. This indicates that letting LLMs see more examples can improve their fallacy understanding ability, but the number of examples must be large enough because we have also seen that when only one shot example is added, the performance of LLMs tends to decline compared to the zero-shot cases.
@@
Table 3: Human evaluation and automated evaluation results on the explanation task. Note that we multiply the human results by 2 to normalize their range to be the same as the GPT-4 results’ range. The reported correlations are Spearman’s rank correlation coefficients. All correlations are extremely significant with $p<0.01$.
Input Question 1 有个问题,鸡蛋干加水会变成鸡蛋吗?
@@
Table 4: Representative case study. We mark correct/partially correct/wrong explanations in green/orange/red.
@@ 3.4 Human Evaluation of Explanation
To verify the effectiveness of our designed automatic GPT-4’s evaluation for Task 3, we randomly select 50 data samples from FLUB, along with outputs from 5 high-performing LLMs for human evaluation by our contracted annotators. From the human evaluation results in Table 3, we observe that:
@@
    Both automatic and human evaluations exhibit a broadly consistent ranking across the selected five models. The GPT-4-Turbo achieves superior performance over all other models. In contrast, human annotators perceive marginal performance disparities among ERNIE-Bot-4.0, Qwen-72B-Chat, and Yi-34B-Chat models. In addition, a notable discrepancy emerges in the evaluation of Qwen-72B-Chat model, where human annotators assign lower ratings than those derived from GPT-4’s automatic evaluation.
@@
    From the results of human evaluation, except for GPT-4-Turbo, which can exceed the passing score of 6, the performance of other LLMs is still not ideal, which shows that the community still needs to further study how to improve the fallacy understanding ability of LLMs.
@@ 3.5 Case Study
To reflect FLUB’s challenge to LLMs, we conduct a case study on the two advanced models with better overall performance in the question explanation task in Table 4. From the first case, we see that GPT-4-Turbo gives a relatively perfect explanation, while ERNIE-Bot-4.0’s answer does not explain the causal relationship clearly although its final conclusion is correct. According to ERNIE-Bot-4.0’s explanation, if the egg is added with water, it can be restored. This is obviously wrong. In the second case which is more difficult, both ERNIE-Bot-4.0 and GPT-4-Turbo easily fail when facing this cunning question. Specifically, ERNIE-Bot-4.0 follows the trap of the question, not clearly stating that “putting heads on the shore” is an impossible operation, but giving a dumbfounding explanation. In comparison, GPT-4-Turbo’s performance is slightly better, but it does not perceive the trap in the question at all, resulting in an answer that is not what is questioned. It can be seen from these two cases that LLMs’ ability to handle cunning questions is still insufficient.
@@ 4 Related Work
4.1 Reasoning Evaluation of LLMs
Our FLUB is for evaluating the fallacy understanding ability of LLMs, which is closely related to the reasoning of LLMs Dong et al. (2023); Chang et al. (2023); Guo et al. (2023). Therefore, this section first reviews related works on the evaluation of commonsense and logical reasoning of LLMs:
Commonsense Reasoning
Existing commonsense reasoning benchmarks include CommonsenseQA Talmor et al. (2019), PIQA Bisk et al. (2020), Social IQA Sap et al. (2019), HellaSWAG Zellers et al. (2019), and MCTACO Zhou et al. (2019). Their task is presented in the form of multiple-choice questions. The recent LLMs reasoning evaluation works Bang et al. (2023); Bian et al. (2023) have demonstrated that LLMs represented by ChatGPT often cannot accurately utilize commonsense knowledge for the reasoning process, resulting in the commonsense reasoning task still being challenging for LLMs.
Logical Reasoning
For logical reasoning data resources, they can be mainly divided into two categories: Natural Language Inference Saha et al. (2020); Tian et al. (2021); Liu et al. (2021) and Multiple-Choice Reading Comprehension Liu et al. (2020, 2022); Wang et al. (2022); Liu et al. (2023a). So far, there have been studies that have conducted in-depth analyses of the performance of LLMs on these two types of tasks. Liu et al. (2023b) show that logical reasoning is very challenging for LLMs, especially for out-of-distribution data samples.
In summary, research on the reasoning ability of LLMs is the focus of current and future LLMs-centric research. The fallacy understanding ability and cunning questions we are concerned about are actually comprehensive challenges of LLMs.
4.2 Humor in NLP
We have noticed that some samples in FLUB contain humorous expressions. Therefore, NLP research on humor Anjum and Lieberum (2023) is instructive for future exploration on FLUB. Particularly, as a representative task in humor research, word game tasks with puns as the core have been continuously paid attention to by researchers Hempelmann (2008); Li et al. (2022); Chen and Soo (2018); Popova and Dadić (2023). According to our statistics, a large proportion of FLUB are cunning questions belonging to word games. Therefore, we believe that how to improve the humor recognition and processing capabilities of LLMs is also the key to improving the performance of LLMs on FLUB.
@@ 5 Conclusion
In this work, we construct FLUB, a high-quality benchmark consisting of cunning questions designed to evaluate the fallacy understanding ability of LLMs. Furthermore, we evaluate advanced LLMs on FLUB. Detailed analyses indicate FLUB is very challenging and of great research value. To date, most existing LLMs still can not understand the fallacy well, which results in them being far from dealing with complex problems in the real world as easily as humans. We believe that the benchmark and the research direction we provide are valuable for the LLMs community.
Limitations
One limitation of FLUB may be that it consists of Chinese data. In particular, many of the cunning questions in FLUB have certain Chinese cultural and language characteristics as backgrounds, which places extremely high demands on LLMs’ knowledge storage. However, as a community that cannot be ignored in the field of NLP, the development of Chinese NLP and Chinese LLMs has been devoted by generations of researchers.
In addition, we are also actively looking for resources to build the English version of FLUB, namely FLUB2.0. But our own resources are limited after all, so we hope that the introduction of FLUB can attract more researchers in the community to pay attention to the importance of fallacy understanding in LLMs, and join the research dedicated to improving the fallacy understanding ability of LLMs.
Ethics Statement
In this paper, we present a new benchmark, FLUB. We have described the details of the collection, preprocessing, and annotation of FLUB. And we ensure that no infringement or unethical behavior occurred during the dataset construction. In terms of the data itself, to ensure that the dataset we need to release in the future meets ethical requirements, we spend lots of energy on data anonymization, data desensitization, improper data cleaning, etc. Besides, the cunning questions we are concerned about come from daily life and are very common. Therefore, the new research direction and tasks we propose will not cause harm to human society.
Appendix A Details of FLUB Annotation
Since the annotation difficulty of different information is different, the salary we pay to the annotators we employ is also different. Specifically, we pay each person who annotates the question type \$0.5 per sample, each person who writes the correct explanation \$1 per sample, and each person who writes the wrong candidates \$2 per sample. It is worth noting that in addition to the junior annotators providing the initial annotation results, we also set three senior annotators with a salary of \$2 per sample, who are responsible for carefully checking the correctness of the annotation results provided by the junior annotators.
Appendix B Translations of Prompts set in FLUB

### v2

@@ 3.1 Experimental Settings
To better reflect the evaluation of FLUB’s fallacy understanding ability of LLMs, we select some advanced LLMs that are widely used in the Chinese community: (1) ERNIE-Bot [15] is a series of closed-sourced commercial LLMs released by Baidu. We evaluate the three latest chat models, including ERNIE-Bot-3.5, ERNIE-Bot-3.5-Turbo, and ERNIE-Bot-4.0. (2) ChatGPT [16] ChatGPT is undoubtedly the hottest model developed by OpenAI. We evaluate GPT-3.5-Turbo and GPT-4-Turbo. (3) ChatGLM3 [17] is the latest open-sourced model of the ChatGLM which is a series of bilingual LLMs. We evaluate the only open-sourced parameter size of ChatGLM3-6B. (4) Qwen [18] is the open-sourced LLMs developed by the Alibaba Group. We select three chat Qwen models, including Qwen-7B-Chat, Qwen-14B-Chat, and Qwen-72B-Chat. (5) Yi [19] series models are open-sourced LLMs trained from scratch by 01-AI. In our experiments, we select Yi-6B-Chat and Yi-34B-Chat to be evaluated on FLUB. (6) Baichuan2 [20] has achieved the competitive performance of its size on many Chinese benchmarks. We select Baichuan2-7B-Chat and Baichuan2-13B-Chat.
When running LLMs inference, for closed-sourced LLMs, we access corresponding models via the official APIs. Meanwhile, open-sourced models are deployed on 1 to 4 NVIDIA A100 GPUs depending on their parameter size.
Table 1: We bold the optimal and underline the suboptimal of closed/open-source models. We report the overall performance by calculating the geometric mean of the three tasks. We color the result that Chain-of-Thought (CoT) brings positive/negative gain as green/red.
Models Open Selection Classification Explanation Overall
Accuracy F-1 Score GPT-4 Score Performance
Source w/o CoT CoT w/o CoT CoT w/o CoT CoT w/o CoT CoT
ERNIE-Bot-3.5-Turbo [15] 32.97 34.65 1.99 6.09 5.78 5.83 7.24 10.72
ERNIE-Bot-3.5 [15] 52.76 38.37 10.33 11.15 6.35 6.22 15.13 13.86
ERNIE-Bot-4.0 [15] 75.66 71.34 11.84 14.42 7.73 8.11 19.06 20.28
GPT-3.5-Turbo [16] 50.48 48.08 3.09 6.15 6.23 7.00 9.91 12.74
GPT-4-Turbo [16] 79.38 82.73 12.31 13.97 8.95 9.21 20.60 22.00
ChatGLM3-6B [17] 35.85 35.01 7.48 9.34 4.98 4.82 11.01 11.64
Qwen-7B-Chat [18] 38.49 33.69 8.00 10.97 5.39 5.65 11.84 11.98
Qwen-14B-Chat [18] 42.57 43.05 10.34 10.44 5.24 6.24 13.21 14.10
Qwen-72B-Chat [18] 58.63 61.51 9.32 12.26 7.34 7.90 15.89 18.13
Yi-6B-Chat [19] 32.37 29.26 8.87 9.84 5.73 5.39 11.81 11.58
Yi-34B-Chat [19] 47.96 48.80 4.74 11.70 6.97 7.52 11.66 16.17
Baichuan2-7B-Chat [20] 43.17 37.17 1.02 4.45 5.48 4.85 6.23 9.29
Baichuan2-13B-Chat [20] 37.05 38.01 3.52 4.58 5.79 5.84 9.11 10.06
Random - 25.00 7.90 - -
Human - 93.35 63.69 - -
@@ 3.2 Automatic Evaluation Results
The main results are presented in Table 1 and we have the following insights:
@@
    For the difficulty of different tasks, the Answer Selection task is the simplest, which shows that LLMs should have a certain ability to distinguish right from wrong when seeing correct and wrong answers. However, we also see that the performance of all models on the Cunning Type Classification task is unsatisfactory, with F-1 scores below 15.0, and some models even perform below random performance. This deficiency may stem from the models’ limited capability to comprehend the semantics of various cunning types.
@@
    For the connection between different tasks, the comparative outcomes among different models across the three tasks are not consistent. Nevertheless, models that exhibit superior performance in the Answer Selection task tend to generate more plausible explanations. This phenomenon reminds us that there is a close relationship between the Answer Selection task and the Fallacy Explanation task. The interaction between these two tasks is very critical for improving the fallacy understanding ability of LLMs.
@@
    For the model performance of different scale parameters, overall, models of larger scale are better equipped to understand cunning texts, which aligns with intuitive expectations. Of course, there are exceptions. We find that for the Qwen and Yi models, as the parameter size increases, the performance of the Cunning Type Classification task decreases. This is because this task requires a deep understanding of the Chinese language, especially the popular Internet language, and we observe that as the Qwen and Yi models become larger, their ability to understand special Internet language becomes poorer.
@@
    For the impact of Chain-of-Thought, to our surprise, Chain-of-Thought (CoT) does not bring stable improvements to LLMs’ fallacy understanding ability. Especially for the Answer Selection and Fallacy Explanation tasks, CoT even has negative impacts on some models. We think there are two main reasons for this phenomenon: (1) We notice that when the model size exceeds 10B, CoT still has positive effects on these two tasks. This reflects the challenge of our tasks, which makes CoT unable to stimulate the small models to have sufficient capabilities to cope with them. (2) For traditional QA tasks (such as commonsense reasoning, mathematical reasoning, etc.), CoT can improve performance because these tasks themselves are relatively logical, and the process of solving their questions can be modeled as the logical reasoning process. Unlike these tasks, our tasks are not very logical problems but require more intuition about the language. Hence, adding intermediate steps by the CoT has no significant effect on our tasks. In summary, our proposed tasks deserve further research to improve the fallacy understanding ability of LLMs.
  5. 5.
    For the overall performance, considering that the performance values of the three sub-tasks are very different, we use the geometric mean to balance the impact of each sub-task and avoid the excessive impact of a single extreme value on the overall performance. We see that the overall performance of each model is basically consistent with common sense, that is, the larger the model, the better the performance, and CoT also brings positive effects. This shows that FLUB is of high quality and suitable to measure the fallacy understanding ability of LLMs from an overall perspective.
  6. 6.
    For the human performance, we see that humans perform well on the Answer Selection and Cunning Type Classification tasks, which reflects the considerable gap in fallacy understanding between human intelligence and LLMs. It also shows that our proposed new benchmark and tasks are conducive to further promoting the progress of LLMs. Note that the reason why the Fallacy Explanation task is not suitable for evaluating human performance is that its automatic evaluation indicator is the GPT-4 Score. We think that using GPT-4 to evaluate explanations written by humans is unreasonable and unnecessary.
@@ 3.3 The Impact of In-context Learning
We select 5 high-performing LLMs to study the impact of in-context learning on LLMs’ fallacy understanding ability. Demonstrations used for in-context learning are randomly selected. As shown in Figure 3, unlike Chain-of-Thought which has no stable positive effect, the LLMs’ performance with in-context learning is basically on the rise as demonstrations increase. This indicates that letting LLMs see more examples can improve their fallacy understanding ability, but the number of examples must be large enough because we have also seen that when only one shot example is added, the performance of LLMs sometimes declines compared to the zero-shot cases.
Table 2: Human evaluation on the explanation task. Note that we multiply the human results by 2 to normalize their range to be the same as the GPT-4 results’ range. The reported correlations are Spearman’s rank correlation coefficients. All correlations are extremely significant with $p<0.01$.
@@
Table 3: We mark correct/partially correct/wrong explanations in green/orange/red.
Input Text 1 {CJK*}UTF8gbsn有个问题,鸡蛋干加水会变成鸡蛋吗?
@@ 3.4 Human Evaluation of Explanation
To verify the effectiveness of our designed automatic GPT-4 score for Task 3, we randomly select 50 data samples from FLUB, along with outputs from 5 high-performing LLMs for human evaluation by our contracted annotators. From the human evaluation results in Table 2, we observe that:
@@
    Both automatic and human evaluations exhibit a broadly consistent ranking across the selected five models. The GPT-4-Turbo achieves superior performance over all other models. In contrast, human annotators perceive marginal performance disparities among ERNIE-Bot-4.0, Qwen-72B-Chat, and Yi-34B-Chat models.
@@
    From the human evaluation results, except for GPT-4-Turbo, which can exceed the passing score of 6, the performance of other LLMs is still not ideal, which shows that the community still needs to further study how to improve the fallacy understanding ability of LLMs.
@@ 3.5 Case Study
To analyze FLUB’s challenges, we conduct case studies on the two advanced models with better performance in the fallacy explanation task in Table 3. From the first case, we see that GPT-4-Turbo gives a relatively perfect explanation, while ERNIE-Bot-4.0’s answer does not explain the causal relationship clearly although its final conclusion is correct. According to ERNIE-Bot-4.0’s explanation, if the egg is added with water, it can be restored. This is obviously wrong. In the second case which is more difficult, both ERNIE-Bot-4.0 and GPT-4-Turbo easily fail when facing these cunning texts. Specifically, ERNIE-Bot-4.0 follows the trap of the input text, not clearly stating that “putting heads on the shore” is an impossible operation, but giving a dumbfounding explanation. In comparison, GPT-4-Turbo’s performance is slightly better, but it does not perceive the trap in the input text at all, resulting in an answer that is not what is questioned. It can be seen from these two cases that LLMs’ ability to handle cunning texts is still insufficient.
@@ 4 Related Work
Reasoning Evaluation of LLMs
Our FLUB is for evaluating the fallacy understanding ability of LLMs, which is closely related to the reasoning of LLMs [21, 22, 23]. Therefore, we first review related works on the commonsense and logical reasoning of LLMs. Commonsense Reasoning: Existing commonsense reasoning benchmarks include CommonsenseQA [24], PIQA [25], Social IQA [26], HellaSWAG [27], and MCTACO [28]. Their task is presented in the form of multiple-choice questions. The recent LLMs reasoning evaluation works [29, 30] have demonstrated that LLMs represented by ChatGPT often cannot accurately utilize commonsense knowledge for the reasoning process. Logical Reasoning: For logical reasoning data resources, they can be mainly divided into two categories: Natural Language Inference [31, 32, 33] and Multiple-Choice Reading Comprehension [34, 35, 36, 37]. [38] show that logical reasoning is very challenging for LLMs, especially for out-of-distribution data samples. In summary, research on reasoning ability is the focus of the LLMs-centric research.
Humor in NLP
We notice that some samples in FLUB contain humorous expressions. Therefore, NLP research on humor [39, 3] is instructive for future exploration on FLUB. Particularly, as a representative humor task, the word game task with puns as the core has been continuously paid attention to by researchers [40, 41, 42, 43]. According to our statistics, a large proportion of FLUB are cunning texts belonging to word games. Therefore, we believe that how to improve the humor recognition and processing capabilities of LLMs is also the key to improving the performance of LLMs on FLUB.
5 Limitations
One limitation of FLUB may be that it consists of Chinese data. In particular, many of the cunning texts in FLUB have certain Chinese cultural and language characteristics as backgrounds, which places extremely high demands on LLMs’ knowledge storage. However, as a community that cannot be ignored in NLP, the development of Chinese LLMs has been devoted by generations of researchers.
6 Ethics Statement
In this paper, we present a new benchmark, FLUB. We have described the details of the collection, preprocessing, and annotation of FLUB. And we ensure that no infringement or unethical behavior occurred during the dataset construction. In terms of the data itself, to ensure that the dataset we need to release in the future meets ethical requirements, we spend lots of energy on data anonymization, data desensitization, improper data cleaning, etc. Besides, the cunning texts we are concerned about come from daily life and are very common. Therefore, the new research direction and tasks we propose will not cause harm to human society.
@@ 7 Conclusion
In this work, we construct FLUB, a high-quality benchmark consisting of cunning texts designed to evaluate the fallacy understanding ability of LLMs. Furthermore, we evaluate advanced LLMs on FLUB. Detailed analyses indicate FLUB is very challenging and of great research value. To date, most existing LLMs still can not understand the fallacy well, which results in them being far from dealing with complex problems in the real world as easily as humans. We believe that the benchmark and the research direction we provide are valuable for the LLMs community.
Acknowledgments and Disclosure of Funding
This research is supported by National Natural Science Foundation of China (Grant No. 62276154), the Natural Science Foundation of Guangdong Province (Grant No. 2023A1515012914), Shenzhen Science and Technology Program (Grant No. WDZC20231128091437002), Basic Research Fund of Shenzhen City (Grant No. JCYJ20210324120012033 and GJHZ202402183000101), the Major Key Project of PCL for Experiments and Applications (PCL2021A06), and Overseas Cooperation Research Fund of Tsinghua Shenzhen International Graduate School (HW2021008).
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
  • Did you include the license to the code and datasets? [Yes] See Section LABEL:gen_inst.
  • Did you include the license to the code and datasets? [No] The code and the data are proprietary.
  • Did you include the license to the code and datasets? [N/A]
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Abstract and Section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 6.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Section 6.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Abstract, Section 3, Appendix A, Appendix B.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Section 2.3 and Section 3.3.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 3.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 2.1 and Section 3.1.
    2. (b)
      Did you mention the license of the assets? [Yes] See Section 2.1.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See Abstract.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section 2.1.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section 2.1.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix C.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Section 2.1.
Figure 4: Our designed prompts without the Chain-of-Thought idea. Task 3(a) is for the texts that are not expressed in the form of inquiries. Task 3(b) is for inquiries.
Figure 5: Our designed prompts with the Chain-of-Thought idea. Task 3(a) is for the texts that are not expressed in the form of inquiries. Task 3(b) is for inquiries.
Figure 6: The designed prompt of automatic evaluation for Task 3.
Appendix A Our Designed Prompts for FLUB